Bank of America has bonds that pay a coupon interest rate of 5.5 percent and mature in 5 years. If an investor has a required rate of return of 7.9 ​percent, what should she be willing to pay for the​ bond? What happens if she pays more or​ less? a. The price she would be willing to pay for the bond is
Answer #1

Answer:

Assume that Face Value of the Bond= 1000

Therefore Coupons payment = 1000*5.5% = 55

Maturity Value = 1000

Calculating Fv= 1000; PMT=55; Time= 5; I/Y= 7.9%,

We get PV= 903.92 (The price she would be willing to pay for the bond)